# BD-14 3065b (TOI-4987b): from giant planet to brown dwarf: evidence for deuterium burning in old age? Key Words.: giant planets – brown dwarfs – techniques: photometric – techniques: spectroscopic – techniques: radial velocities1 Introduction

### v1

@@ 3.3 Frequency Analysis
We performed a frequency analysis to detect all signals in RVs. Our observations had a baseline of around 500 days, providing a frequency resolution of approximately . We observed a non-linear long trend in the dataset, with an unconstrained period, as seen in the periodogram. After fitting the trend, we identified a significant period of 4.3 days and its one-day alias, consistent with the transits from TESS photometry (see Fig. 13).
@@
3.4 Joint RVs and transits modelling
@@
The process of modeling transits and radial velocities was carried out simultaneously using the allesfitter package. The fitting routine, as described in Günther & Daylan (2021), is based on several packages, including ellc for light-curve and RV models (Maxted, 2016), aflare for flare modeling (Davenport et al., 2014), dynesty for static and dynamic Nested Sampling (Speagle, 2020), emcee for MCMC sampling (Foreman-Mackey et al., 2013), and celerite for GP models (Foreman-Mackey et al., 2017). The allesfitter package offers a range of orbital and transit/eclipse models, accommodating multiple exoplanets, multi-star systems, transit-timing variations, phase curves, stellar variability, starspots, stellar flares, and various systematic noise models, including Gaussian processes.
@@
To obtain the optimal conjunction time, period, inclination, eccentricity, and argument of periastron, which were later used as priors for joint modeling, we performed a fit of photometry alone, as the secondary eclipse is visible in TESS photometry (see Fig. 14). The LCs were already corrected for variability using the procedure outlined in Section 2.1, and we employed the quadratic limb darkening law with coefficients q1 and q2. To ensure accuracy and avoid degeneracy between these coefficients and transit depth, we adopted Gaussian priors within around the values from Claret (2017), which presents limb-darkening coefficients specifically computed for the TESS photometry for a wide range of effective temperatures, gravities, metallicities, and microturbulent velocities. Although modeling the pulsations simultaneously with the transits may be preferred, it will not impact the results since pulsations are much slower than the transit duration. We conducted two separate fits based on the stellar parameters and dilution factors obtained from two-star and three-star SED modeling. Table 4 shows the priors and posteriors of fitted parameters, while Table 5 presents the transit parameters and physical parameters derived from Tables 3 and 4, respectively. The results are plotted in Fig. 14.
@@
The stellar radius is an important parameter that leads to the companion radius computed through the transit depth. We need to double-check that the radius of BD-14 3065b is correct, as the discussion of this work is built upon its value. As a first check, we modeled the 2-star SED, assuming that the star found in the high-resolution imaging is an unassociated companion with its own age and distance. However, such a model led to similar parameters of the primary star ( M, a radius of R). As a second check, to ensure the reliability of reported stellar spectral parameters, which are crucial in our SED modeling, we conducted a search for values in the relevant literature. The Sixth Data Release of the Radial Velocity Experiment (RAVE) reported stellar parameters of K and . Gaia reported stellar parameters of K and using the photometry333https://gea.esac.esa.int/archive/documentation/GDR3/Data_analysis/chap_cu8par/sec_cu8par_apsis/ssec_cu8par_apsis_gspphot.html. These values are consistent with our own findings. Additionally, it is important to note that unresolved binarity is not expected to impact these values significantly (El-Badry et al., 2018). We are, therefore, confident that the reported stellar parameters are reliable. Finally, the impact of an unresolved binary star on the radius of a transiting planet has been the subject of much discussion in the literature. The ratio of the planet’s true radius , to the observed radius when assuming a single star with no companions, can be described as:
@@ 4.2 Deuterium burning
In order to investigate the possibility of overheating in BD-14 3065b, we derived its thermal emission through secondary eclipse depth analysis. The depth is proportional to the square of the radius ratio and the temperature ratio of the host and transiting companion. We analyzed the TESS LCs and found a secondary eclipse depth of parts per thousand (ppt). The secondary eclipse fit can be observed in Fig. 14. It is important to note that the TESS bandpass window444https://archive.stsci.edu/files/live/sites/mast/files/home/missions-and-data/active-missions/tess/_documents/TESS_Instrument_Handbook_v0.1.pdf is centered at roughly 800 nm and spans nm, which means that the relative brightness of the planet to the star is only measured at these wavelengths. We used this relative brightness to determine the effective temperature of BD-14 3065b, employing the equation used for the secondary eclipse depth from Charbonneau et al. (2005):
where is the TESS spectral response function and is the flux of companion or star. In order to derive BD-14 3065b’s effective temperature, we simplify the equation by assuming blackbody profiles to describe fluxes in the TESS bandpass window and zero geometric albedo. Our findings indicate that BD-14 3065b has an effective blackbody temperature of . This temperature is notably higher than the equilibrium temperature that would result from star heating alone. Consequently, the secondary eclipse has revealed that the planet is experiencing overheating, which would increase the deuterium burning rate. The validity of this conclusion is contingent upon the assumption of BD-14 3065b’s blackbody spectrum. However, it remains uncertain whether the eclipse spectra of the hottest atmospheres align with blackbody curves. Studies of individual planets with HST and Spitzer have provided observational evidence in support of proximity to blackbody (e.g., Kreidberg et al., 2018; Mansfield et al., 2018; Arcangeli et al., 2019). Nevertheless, recent research has highlighted a rise in brightness for wavelengths shorter than 1.3 m due to the combined opacities of H, TiO, VO, and FeH (e.g., Changeat et al., 2022; Coulombe et al., 2023). On the other hand, according to observations from Coulombe et al. (2023), TESS observation of planet thermal emission aligns well with the black body spectrum. Consequently, we are of the opinion that our assumption is reasonable.
@@ 4.3 Overheating of BD-14 3065b
It is a well-established fact that the increase in entropy due to the interior heating of a planet leads to an increase in its radius (Arras & Bildsten, 2006; Marleau & Cumming, 2014). In this regard, we aim to investigate the possibility of BD-14 3065b being heated due to the interaction and dissipation between the star and the planet. Analysis of the TESS photometry data reveals that the stellar spin frequency is in sync with the planet’s orbital mean motion, indicating that the synchronization timescale is comparable to, or smaller than, the star’s age. The synchronization timescale for the planet’s spin angular frequency is smaller than the orbital synchronization timescale (Bodenheimer et al., 2001). Therefore, the synchronization process of planet rotation cannot contribute significantly to the heating of the planetary interior.
Another potential heating source is the dissipation of a residual eccentricity, as indicated by a slightly eccentric orbit observed from a secondary eclipse. The rate of eccentric orbital energy dissipation inside the planet is proportional to the value of eccentricity, and this rate would decrease over time as the orbit becomes more circularized. The rate also depends on the circularization timescale or, more specifically, on the planet’s and star’s tidal quality factors, and while the circularization timescale is several orders smaller as stars become sub-giant, most of the eccentric orbital energy is dissipated inside the star rather than the planet. Therefore, while the planet may indeed be heated from a residual eccentricity, this source of heating weakens over time as the orbit becomes less eccentric. As this heating source was stronger in the past and affects many eccentric systems in literature, there is no obvious connection to BD-14 3065’s large radius. To calculate the rate of dissipation of the planet’s eccentric orbital energy within the planet’s interior, we used an equation from Bodenheimer et al. (2001). Assuming the planet’s tidal quality factor , we computed the value of erg s.
@@
To compare the rate of energy transfer to the planet’s interior to the one coming from an eccentric orbit and determine if the observed radius of BD-14 3065b could be matched for a realistic fraction of stellar irradiation transported into the planet’s interior, we used the hierarchical Bayesian code developed by Sarkis et al. (2021). The code utilizes the planetary evolution model completo21 (Mordasini et al., 2012) and fully non-gray atmospheric models from the petitCODE (Mollière et al., 2015) to connect the interior structure models to the observed physical properties such as the planet’s radius and mass, and the stellar luminosity and semimajor axis. The code assumes the planet is in a steady state and thus does not calculate the planetary thermal evolution. The planet models assume no central core, and the heavy elements (made up entirely of water) are equally distributed throughout the interior. A study by Sarkis et al. (2021) found that the heating efficiency increases with increasing equilibrium temperature and reaches a maximum of 2.5% at around 1860 K, beyond which the efficiency decreases. When assuming the highest heating efficiency of 2.5%, we computed the value of erg s, which is an order of magnitude higher than the energy dissipated from an eccentric orbit. However, this amount of energy can produce a radius only of about , assuming the stellar fraction of heavy element mass. Even if we assume a heating efficiency of 20%, we can barely reach . There is no known process that can transfer such a high fraction of stellar irradiation into the interior of the planet. Therefore, we conclude that while several anticipated processes may increase the internal temperature of the planet, they would need to be unreasonably efficient to account for the observed radius. However, such processes would also significantly increase the rate of deuterium reactions. This appears to be the most plausible scenario. Alternatively, we can speculate that the planet may have experienced a recent collision that increased its internal temperature.
@@ 4.5 Follow-up prospects
BD-14 3065b presents itself as a highly promising target for further atmospheric characterization through secondary eclipse observations. The planet’s equilibrium temperature classifies it as an ultra-hot object that is comparable to WASP-18b, whose thermal emission spectrum was previously observed using the James Webb Space Telescope (Coulombe et al., 2023). The most prominent features that can be observed in such an atmosphere include HO, CO, FeH, TiO, and VO absorption. A simple atmospheric model of the anticipated thermal emission spectrum in the NIRSpec and NIRISS wavelength ranges was created (see Fig. 19) using the picaso code (Batalha et al., 2019) alongside the Sonora brown dwarf evolutionary models (Marley et al., 2021). In a similar fashion to WASP-18b, the James Webb Space Telescope would be able to determine the metallicity and carbon-to-oxygen (C/O) ratio of BD-14 3065b.
@@
C/O ratios have been measured for several transiting planets (e.g., Madhusudhan et al., 2011; Moses et al., 2013; Line et al., 2014; Arcangeli et al., 2018; Ahrer et al., 2023). Beatty et al. (2018) conducted a study on the brown dwarf CWW 89Ab transiting a 3 Gyr old Sun-like star. The team used the Spitzer/IRAC to observe the secondary eclipses, and they discovered a significant - overluminosity of CWW 89Ab compared to the predicted values from the evolutionary models. The authors discussed the potential causes of this overluminosity, ruling out both stellar irradiation and tidal heating as likely explanations. They suggested that a dayside temperature inversion, resulting from a super-stellar C/O ratio, could be responsible for this phenomenon. If this is indeed the case, it would imply that CWW 89Ab is a object that formed via the core accretion process. The measurement of the C/O ratio would be essential to gain a better understanding of this system. It is also worth noting that, in contrast to BD-14 3065b, CWW 89Ab does not exhibit an inflated radius. We may also mention a recent study by Yan et al. (2023) in which authors used ground-based high-resolution spectrograph CRIRES to observe dayside thermal emission spectra of two ultra-hot Jupiters and detected strong CO emission lines in both planets. BD-14 3065b is an ideal target for such a study.
@@
The C/O ratio is an important characteristic proposed to constrain the formation of giant planets (Öberg et al., 2011). Generally, a stellar C/O ratio is expected for companions that form through gravitational instabilities in a cloud, where all materials mix, and planets that form within the snow lines. A superstellar C/O ratio could indicate that the atmosphere formed from disk gas outside the water snowline, possibly via core accretion. It was later discussed that such simple division is not realistic and that several complex and interconnected processes affect the planetary formation, such as disc elemental composition, chemical evolution of the protoplanetary disc, planetary formation and migration, atmospheric evolution, and others (see Mollière et al. (2022) for more details). It is still a long way off, if even possible at all, to connect atmospheric abundances with formation processes for individual systems. However, we can use the observed atmospheric composition to study the importance of various formation aspects in an attempt to reproduce observations. Furthermore, the atmospheric abundance constraints obtained with new and upcoming instruments will be crucial in developing planet formation models. Instruments such as GRAVITY, CRIRES, and recent and upcoming facilities such as JWST, ARIEL, and the ELT will obtain atmosphere composition for many planets.
@@ 5 Summary
We have presented an analysis of the substellar object BD-14 3065b transiting a subgiant F-type star. BD-14 3065b has an orbital period of P = 4.2889731 ± 0.0000052 days, and the estimated age of the system is about 2 Gyr, making BD-14 3065b an extremely inflated object for its mass and age. This study focused on the two most notable differences between BD-14 3065b and other systems in the literature, attempting to link them to the observed radius. Firstly, BD-14 3065b is located at the deuterium burning boundary, and secondly, its primary star is an evolved subgiant. The conclusion drawn is that the radius could be explained by the enhanced rate of deuterium burning caused by sufficient heating of BD-14 3065b’s interior. Finally, detection of the secondary eclipse with TESS photometry enables a precise determination of the eccentricity and reveals BD-14 3065b has a brightness temperature of K. Hence, BD-14 3065b is considered an excellent target for future atmospheric characterization through secondary eclipse observations. Such an investigation would provide valuable insights into the processes responsible for the planet’s atmospheric heating, as well as its formation history, by comparing its C/O ratio to that of its host star. This would be of significant value, given that BD-14 3065b sits at the giant planet/brown dwarf boundary.

### v2

@@ 3.3 Frequency Analysis
We performed a frequency analysis to detect all signals in RVs. Our observations had a baseline of around 500 days, providing a frequency resolution of approximately . We observed a non-linear long trend in the dataset, with an unconstrained period as seen in the periodogram. This signal corresponds to the close stellar companion, which we accounted for in the three-component SED modeling. The period, eccentricity, and mass of the companion remain unconstrained, although its minimum mass is in the stellar regime (assuming the shortest possible period). After fitting the trend, we identified a significant period of 4.3 days and its one-day alias, consistent with the transits from TESS photometry (see Fig. 13).
@@
3.4 Joint RVs, transits and eclipses modelling
@@
The process of modeling transits, eclipses, and radial velocities was carried out simultaneously using the allesfitter package. The fitting routine, as described in Günther & Daylan (2021), is based on several packages, including ellc for light-curve and RV models (Maxted, 2016), aflare for flare modeling (Davenport et al., 2014), dynesty for static and dynamic Nested Sampling (Speagle, 2020), emcee for MCMC sampling (Foreman-Mackey et al., 2013), and celerite for GP models (Foreman-Mackey et al., 2017). The allesfitter package offers a range of orbital and transit/eclipse models, accommodating multiple exoplanets, multi-star systems, transit-timing variations, phase curves, stellar variability, starspots, stellar flares, and various systematic noise models, including Gaussian processes.
@@
To obtain the optimal conjunction time, period, inclination, eccentricity, and argument of periastron, which were later used as priors for joint modeling, we performed a fit of photometry alone, as the secondary eclipse is visible in TESS photometry (see Fig. 14). The LCs were already corrected for variability using the procedure outlined in Section 2.1, and we employed the quadratic limb darkening law with coefficients q1 and q2. To ensure accuracy and avoid degeneracy between these coefficients and transit depth, we adopted Gaussian priors within around the values from Claret (2017), which presents limb-darkening coefficients specifically computed for the TESS photometry for a wide range of effective temperatures, gravities, metallicities, and microturbulent velocities. Although modeling the pulsations simultaneously with the transits may be preferred, it will not impact the results since pulsations are much slower than the transit duration. We conducted two separate fits based on the stellar parameters and dilution factors obtained from two-star and three-star SED modeling.
We conducted a joint fit of our data, incorporating TRES and Pucheros+ RVs processed in accordance with the procedures outlined in Section 2. In the case of the TRES data, we did not use RVs obtained from the TODCOR analysis, as we discovered that the signal is associated with the F star, rendering any significant difference unlikely. To test this assertion, we modeled the orbit using RVs obtained from the TODCOR analysis in Section 2, and we found that all parameters were consistent within . Additionally, the RVs used in Section 2 were obtained via a more sophisticated method, utilizing the multi-order approach. Table 3 shows the priors and posteriors of fitted parameters, while Table 4 presents the transit parameters and physical parameters derived from Tables 2 and 3, respectively. The results are plotted in Fig. 14.
@@
The stellar radius is an important parameter that leads to the companion radius computed through the transit depth. We need to double-check that the radius of BD-14 3065b is correct, as the discussion of this work is built upon its value. As a first check, we modeled the 2-star SED, assuming that the star found in the high-resolution imaging is an unassociated companion with its own age and distance. However, such a model led to similar parameters of the primary star ( M⊙, a radius of R⊙). As a second check, to ensure the reliability of reported stellar spectral parameters, which are crucial in our SED modeling, we conducted a search for values in the relevant literature. The Sixth Data Release of the Radial Velocity Experiment (RAVE) reported stellar parameters of K and . Gaia reported stellar parameters of K and using the photometry444https://gea.esac.esa.int/archive/documentation/GDR3/Data_analysis/chap_cu8par/sec_cu8par_apsis/ssec_cu8par_apsis_gspphot.html. These values are consistent with our own findings. Additionally, it is important to note that unresolved binarity is not expected to impact these values significantly (El-Badry et al., 2018). We are, therefore, confident that the reported stellar parameters are reliable. Finally, the impact of an unresolved binary star on the radius of a transiting planet has been the subject of much discussion in the literature. The ratio of the planet’s true radius , to the observed radius when assuming a single star with no companions, can be described as:
@@ 4.2 Deuterium burning
In order to investigate the possibility of overheating in BD-14 3065b, we derived its thermal emission through secondary eclipse depth analysis. The depth is proportional to the square of the radius ratio and the temperature ratio of the host and transiting companion. In Section 3.4, we analyzed the TESS LCs and found a secondary eclipse depth of parts per thousand (ppt). The secondary eclipse fit can be observed in Fig. 14. It is important to note that the TESS bandpass window555https://archive.stsci.edu/files/live/sites/mast/files/home/missions-and-data/active-missions/tess/_documents/TESS_Instrument_Handbook_v0.1.pdf is centered at roughly 800 nm and spans nm, which means that the relative brightness of the planet to the star is only measured at these wavelengths. We used this relative brightness to determine the effective temperature of BD-14 3065b, employing the equation used for the secondary eclipse depth from Charbonneau et al. (2005):
where is the TESS spectral response function and is the flux of companion or star. In order to derive BD-14 3065b’s effective temperature, we simplify the equation by assuming blackbody profiles to describe fluxes in the TESS bandpass window and zero geometric albedo. Our findings indicate that BD-14 3065b has an effective blackbody temperature of . This temperature is notably higher than the equilibrium temperature that would result from star heating alone. Consequently, the secondary eclipse has revealed that the planet is experiencing overheating, which would increase the deuterium burning rate. The validity of this conclusion is contingent upon the assumption of BD-14 3065b’s blackbody spectrum. However, it remains uncertain whether the eclipse spectra of the hottest atmospheres align with blackbody curves. Studies of individual planets with HST and Spitzer have provided observational evidence in support of proximity to blackbody (e.g., Kreidberg et al., 2018; Mansfield et al., 2018; Arcangeli et al., 2019). Nevertheless, recent research has highlighted a rise in brightness for wavelengths shorter than 1.3 m due to the combined opacities of H-, TiO, VO, and FeH (e.g., Changeat et al., 2022; Coulombe et al., 2023). On the other hand, according to observations from Coulombe et al. (2023), TESS observation of planet thermal emission aligns well with the black body spectrum. Consequently, we are of the opinion that our assumption is reasonable.
@@ 4.3 Overheating of BD-14 3065b
It is a well-established fact that the increase in entropy due to the interior heating of a planet leads to an increase in its radius (Arras & Bildsten, 2006; Marleau & Cumming, 2014). In this regard, we aim to investigate the possibility of BD-14 3065b being heated due to the interaction and dissipation between the star and the planet. Analysis of the TESS photometry data reveals that the stellar spin frequency is comparable with the planet’s orbital mean motion, indicating that the synchronization timescale is comparable to the star’s age. The synchronization timescale for the planet’s spin angular frequency is smaller than the orbital synchronization timescale (Bodenheimer et al., 2001). Therefore, the synchronization process of planet rotation cannot contribute significantly to the heating of the planetary interior.
Another potential heating source is the dissipation of a residual eccentricity, as indicated by a slightly eccentric orbit observed from a secondary eclipse. The rate of eccentric orbital energy dissipation inside the planet is proportional to the value of eccentricity, and this rate would decrease over time as the orbit becomes more circularized. The rate also depends on the circularization timescale or, more specifically, on the planet’s and star’s tidal quality factors, and while the circularization timescale is several orders smaller as stars become sub-giant, most of the eccentric orbital energy is dissipated inside the star rather than the planet. Therefore, while the planet may indeed be heated from a residual eccentricity, this source of heating weakens over time as the orbit becomes less eccentric. As this heating source was stronger in the past and affects many eccentric systems in literature, there is no obvious connection to BD-14 3065’s large radius. To calculate the rate of dissipation of the planet’s eccentric orbital energy within the planet’s interior, we used an equation from Bodenheimer et al. (2001). Assuming the planet’s tidal quality factor , we computed the value of erg s-1.
@@
To compare the rate of energy transfer to the planet’s interior to the one coming from an eccentric orbit and determine if the observed radius of BD-14 3065b could be matched for a realistic fraction of stellar irradiation transported into the planet’s interior, we used the hierarchical Bayesian code developed by Sarkis et al. (2021). The code utilizes the planetary evolution model completo21 (Mordasini et al., 2012) and fully non-gray atmospheric models from the petitCODE (Mollière et al., 2015) to connect the interior structure models to the observed physical properties such as the planet’s radius and mass, and the stellar luminosity and semimajor axis. The code assumes the planet is in a steady state and thus does not calculate the planetary thermal evolution. The planet models assume no central core, and the heavy elements (made up entirely of water) are equally distributed throughout the interior. A study by Sarkis et al. (2021) found that the heating efficiency increases with increasing equilibrium temperature and reaches a maximum of 2.5% at around 1860 K, beyond which the efficiency decreases. When assuming the highest heating efficiency of 2.5%, we computed the value of erg s-1, which is an order of magnitude higher than the energy dissipated from an eccentric orbit. However, this amount of energy can produce a radius only of about , assuming the stellar fraction of heavy element mass. Even if we assume a heating efficiency of 20%, we can barely reach . There is no known process that can transfer such a high fraction of stellar irradiation into the interior of the planet. Therefore, we conclude that while several anticipated processes may increase the internal temperature of the planet, they would need to be unreasonably efficient to account for the observed radius. However, such processes would also significantly increase the rate of deuterium reactions. This appears to be the most plausible scenario. Alternatively, we can speculate that the planet may have experienced a recent collision that increased its internal temperature.
@@ 4.5 Follow-up prospects
BD-14 3065b presents itself as a highly promising target for further atmospheric characterization through secondary eclipse observations. The planet’s equilibrium temperature classifies it as an ultra-hot object that is comparable to WASP-18b, whose thermal emission spectrum was previously observed using the James Webb Space Telescope (Coulombe et al., 2023). The most prominent features that can be observed in such an atmosphere include H2O, CO, FeH, TiO, and VO absorption. A simple atmospheric model of the anticipated thermal emission spectrum in the NIRSpec and NIRISS wavelength ranges was created (see Fig. 19) using the picaso code (Batalha et al., 2019) alongside the Sonora brown dwarf evolutionary models (Marley et al., 2021). In a similar fashion to WASP-18b, the James Webb Space Telescope would be able to determine the metallicity and carbon-to-oxygen (C/O) ratio of BD-14 3065b.
@@
C/O ratios have been measured for several transiting planets (e.g., Madhusudhan et al., 2011; Moses et al., 2013; Line et al., 2014; Arcangeli et al., 2018; Ahrer et al., 2023). Beatty et al. (2018) conducted a study on the brown dwarf CWW 89Ab transiting a 3 Gyr old Sun-like star. The team used the Spitzer/IRAC to observe the secondary eclipses, and they discovered a significant - overluminosity of CWW 89Ab compared to the predicted values from the evolutionary models. The authors discussed the potential causes of this overluminosity, ruling out both stellar irradiation and tidal heating as likely explanations. They suggested that a dayside temperature inversion, resulting from a super-stellar C/O ratio, could be responsible for this phenomenon. If this is indeed the case, it would imply that CWW 89Ab is a object that formed via the core accretion process. The measurement of the C/O ratio would be essential to gain a better understanding of this system. It is also worth noting that, in contrast to BD-14 3065b, CWW 89Ab does not exhibit an inflated radius. We may also mention a recent study by Yan et al. (2023) in which authors used ground-based high-resolution spectrograph CRIRES+ to observe dayside thermal emission spectra of two ultra-hot Jupiters and detected strong CO emission lines in both planets. BD-14 3065b is an ideal target for such a study.
@@
The C/O ratio is an important characteristic proposed to constrain the formation of giant planets (Öberg et al., 2011). Generally, a stellar C/O ratio is expected for companions that form through gravitational instabilities in a cloud, where all materials mix, and planets that form within the snow lines. A superstellar C/O ratio could indicate that the atmosphere formed from disk gas outside the water snowline, possibly via core accretion. It was later discussed that such simple division is not realistic and that several complex and interconnected processes affect the planetary formation, such as disc elemental composition, chemical evolution of the protoplanetary disc, planetary formation and migration, atmospheric evolution, and others (see Mollière et al. (2022) for more details). It is still a long way off, if even possible at all, to connect atmospheric abundances with formation processes for individual systems. However, we can use the observed atmospheric composition to study the importance of various formation aspects in an attempt to reproduce observations. Furthermore, the atmospheric abundance constraints obtained with new and upcoming instruments will be crucial in developing planet formation models. Instruments such as GRAVITY, CRIRES+, and recent and upcoming facilities such as JWST, ARIEL, and the ELT will obtain atmosphere composition for many planets.
@@ 5 Summary
We have presented an analysis of the substellar object BD-14 3065b transiting a subgiant F-type star. BD-14 3065b has an orbital period of P = 4.2889731 ± 0.0000052 days, and the estimated age of the system is about 2 Gyr, making BD-14 3065b, with a mass of and a radius of , an extremely inflated object for its mass and age. This study focused on the two most notable differences between BD-14 3065b and other systems in the literature, attempting to link them to the observed radius. Firstly, BD-14 3065b is located at the deuterium burning boundary, and secondly, its primary star is an evolved subgiant. The conclusion drawn is that the radius could be explained by the enhanced rate of deuterium burning caused by sufficient heating of BD-14 3065b’s interior. Finally, detection of the secondary eclipse with TESS photometry enables a precise determination of the eccentricity and reveals BD-14 3065b has a brightness temperature of K. Hence, BD-14 3065b is considered an excellent target for future atmospheric characterization through secondary eclipse observations. Such an investigation would provide valuable insights into the processes responsible for the planet’s atmospheric heating, as well as its formation history, by comparing its C/O ratio to that of its host star. This would be of significant value, given that BD-14 3065b sits at the giant planet/brown dwarf boundary.